# Efficient Swept Volume-Based Trajectory Generation for Arbitrary-Shaped Ground Robot Navigation

### v1

@@ I INTRODUCTION
Safe navigation of arbitrary-shaped ground robots in constrained environments is an indispensable requirement for real-world autonomy in many scenarios. For example, forklift trucks transporting long timber in the warehouse demand precise whole-body motion planning to avoid collisions in confined aisles. Traditional methods [1, 2, 3] that conservatively approximate robots with convex shapes largely waste the navigable space, rendering them impractical for such applications. Thus, trajectory planners that explicitly model the robot’s true geometry are necessary for efficient and safe deployment in spatially constrained scenes.
Recent frameworks like Robot-Centric Euclid Signed Distance Field (RC-ESDF) [4] and Implicit Swept Volume Signed Distance Field (SVSDF) [5] enable precise geometry-aware robot collision assessment, overcoming the limitations of direct convex approximations. However, RC-ESDF adopts a naive A* searcher that renders a single path candidate, which ignores the possibility of other potentially feasible paths. Furthermore, its discrete sampling collision evaluation in back-end optimization risks missing the sub-resolution narrow collisions in the scene. Unlike RC-ESDF, SVSDF ensures provably accurate Continuous Collision Avoidance (CCA) [6] by leveraging an iterative swept volume approach. However, the SVSDF planner also employs an A*-based searcher, which suffers the same problem as RC-ESDF’s front end. Moreover, SVSDF’s prohibitive runtime makes it impractical for real-world autonomous applications, which primarily stems from its monolithic back-end formulation that attempts to solve the entire whole-body robot trajectory from start to goal in a single step. In practical ground robot navigation, we observe that finding feasible navigation options rarely succeeds with a single front-end path candidate. Besides, optimizing the whole body SE(2) trajectory is necessary only in localized regions with restricted obstacles, while simpler $\mathbb{R}^{2}$ planning suffices elsewhere. However, both RC-ESDF and SVSDF lack a geometry-aware front end that can provide multiple navigable options, identify critical regions, and adaptively activate whole-body optimization only where needed—leading to computational inefficiency.
@@ II RELATED WORKS
In robot motion planning, existing works [7], [8] usually model the robot’s geometry as a mass point and inflate the obstacle based on its radius to construct configuration space for collision avoidance. Liu et al. [9] encodes the robot shape as an ellipsoid to search the safe path for drones. Euclidean Signed Distance Fields (ESDF) and corridor-based methods are widely employed to construct collision-aware cost functions for whole-body motion planning trajectory optimization.
@@
The ESDF serves as a map representation that encodes the distance to the nearest obstacle surface. In this approach, the robot’s geometry is often approximated as a sphere [10], [11] to simplify distance queries. Collision-free motion is guaranteed by ensuring the ESDF value (i.e., the minimum distance to obstacles) at each trajectory sample point exceeds the radius of the corresponding sphere. While computationally efficient, this spherical approximation may lead to overly conservative trajectories, as the spheres only coarsely enclose the robot’s true geometry. In contrast, corridor-based methods construct trajectories by defining collision-free regions as convex geometric primitives, including polyhedra [1], spheres [2], and axis-aligned bounding boxes [3]. Li et al. [12] employs multiple spheres to approximate the robot’s shape, enabling a more flexible and accurate representation of its geometry within the collision-free corridors. These primitives impose linear constraints on the trajectory optimization problem, ensuring the robot’s motion remains within the corridor. To account for the robot’s spatial footprint, the corridor framework often models the robot as an ellipsoid [13], a union of spheres [12], [14] and enforces its containment within the corridor’s convex polyhedra. All the above-mentioned methods share a core principle: generating a convex polytope that encapsulates the robot’s geometry to guarantee collision avoidance. However, ensuring continuous collision-free trajectories while preserving the accurate representation of a robot’s non-convex geometric shape remains a significant challenge.
@@
Inspired by the flexibility of the ESDF in representing arbitrary obstacle geometries, Geng et al. [4] proposed a novel approach that inverts the traditional ESDF paradigm: instead of computing distances from obstacles, the method evaluates distances from the robot’s surface to obstacles. By reformulating the ESDF around the robot’s geometry, the framework actively “pushes” the robot away from obstacles during trajectory optimization. However, this method faces two critical limitations. First, like traditional ESDF, the robot-centric ESDF relies on discrete sampling of the robot’s surface along trajectory which may results in “tunnel effect”. Second, the gradient direction derived from the robot-centric ESDF does not reflect the optimal escape direction for the robot.
@@
To address these issues, SVSDF[6] innovatively integrates the concept of swept volume into trajectory optimization. It represents the robot’s surface as the zero-level set of an implicit SDF and integrates swept volume to model the space-time continuum of its motion to achieve resolution-independent CCA. Additionally, the SVSDF is calculated by finding the radius of the smallest sphere at an obstacle point that is tangent to the swept volume boundary which is always along the optimal gradient for collision avoidaance. However, precisely considering the robot’s actual geometry in a single optimization problem from start to end for the continuous collision-free trajectory is time-consuming. It causes much pressure for the optimizer to converge to a satisfied minimum. Thus, we present an adaptive coarse-to-fine trajectory generation paradigm, accurately representing the robot’s geometry only when it is needed, and using simpler planners elsewhere to speed up the process.
@@ IV Topological Path Generation
To maximize the exploration of various topological paths in a narrow space while maintaining efficiency, in this stage, we use the minimum inflation margin which is the inscribed circumference of the robot as shown in Fig. 3(a), to inflate the occupancy grid map, representing the coarse navigability at this initial stage. After that, we construct topological roadmaps based on [15] with a rough consideration of the robot’s shape.
@@ 6             $p_{new}\leftarrow\neg\textbf{Safe}$;
7             while $\neg\textbf{Safe}(p_{new})$land $N_{attempt}$ $<$ $N_{max}$ do
@@ 11                   $\mathcal{M}_{yaw}\leftarrow\textbf{UpdateState}(\mathcal{\theta}_{new},p_{new},\mathcal{M})$;
In [15], redundant zigzag paths in obstacle-dense regions are shortened while treating the robot as a mass point and ignoring the actual shape. This will result in missing regions that can be passed in a particular orientation. Therefore, we redesigned the shortcut process to generate a refined topologically equivalent shortcut path $P_{s}$ considering the any-shape robot’s physical geometry (refer to Alg. 1 and Fig. 4). Initially, we denote the original detoured path (blue line) found by the depth-first search from the roadmap as $P_{i}$ and the expected final simplified path (green line) as $P_{s}$. Then, the algorithm uniformly discretizes $P_{i}$ into a set of points $P_{d}=\{p_{1},p_{2},...,p_{n}\}$ (blue point in the path) based on the grid map resolution and initializes the simplified path $P_{s}$ with the start point $p_{1}$ (Line 1). For each point $p_{d}$ in $P_{d}$, the algorithm performs visibility checking between $P_{s}.back()$ and $p_{d}$ as in [15] (Line 2). When visibility is blocked by an obstacle, the algorithm identifies the obstruction point as $p_{c}$ (Line 4).
@@
Then, we expect to push the obstructed configuration in $p_{c}$ away from the occupancy to generate new collision-free waypoints. Inspired by [4], to handle the arbitrary-shaped robot, we construct body-frame ESDF inside the robot geometry $\mathcal{M}$ centered at obstruction point $p_{c}$ to render the obstacle-avoiding direction, as shown in Fig. 3(b). The robot’s initial orientation at the obstruction point $p_{c}$ (red T-shape) is important to construct the appropriate body-frame ESDF for guiding collision point adjustment. It is determined by applying Cubic Spline Interpolation between the initial and final orientation at $P_{s}.back()$ and $p_{d}$ (blue T-shape), respectively, and then storing the resulting orientation with the geometric representation and position in $\mathcal{M}_{yaw}$ for generating the ESDF (Line 5). we build an Axis-Aligned Bounding Box (AABB) to extract all nearby obstacle points $X_{obs}$ relative to $p_{c}$ (Line 8). The SDF value at any position $x_{obs}$ in $X_{obs}$ is defined as the minimum distance from it to the nearest robot surface, with positive values indicating the robot’s exterior and negative values indicating its interior, which can be efficiently computed via a generalized winding number method [16] within LIBIGL [^2] https://libigl.github.io/ (Line 9). The gradient in the robot local frame $\nabla\mathcal{SDF}_{robot}^{\mathcal{M}}$ is defined as the direction of the steepest ascent to the nearest robot surface from the query point $x_{obs}$ and computed using central finite differences of distance values between adjacent grid cells. Subsequently, the gradient of the robot SDF in the world frame $\nabla\mathcal{SDF}_{world}^{\mathcal{M}}$ which directs away from the obstacle is obtained with a proper coordinate transform as follows:
@@
where ${R}^{-1}(p_{c})$ is the inverse of the robot’s rotation matrix at $p_{c}$. The norm and direction of the gradient (pink arrow) can push the center of robot ESDF in obstruction point $p_{c}$ away from the obstacle with a magnitude based on the SDF value and minimum safe margin (Line 10). The resultant position after applying SDF-based steering once might remain collided with the confined environments. Thus, we repeat this process within the maximum allowed attempts (Line 7). If the robot collides after reaching the limitation attempts, the last $p_{new}$ along with its orientation $\theta_{new}$ is temporally designated as a new waypoint for further SE(2) refinement (Line 12). The refining process continues until the last point is reached. After that, uniform visibility deformation (UVD) [15] is implemented to distinguish topological equivalence and eliminate redundant paths effectively. Finally, Alg. 1 generates a new shortened and smoother path consisting of a series of waypoints with feasible orientation stored according to the actual geometry of the robot while preserving the topology attribute.
@@ V SE(2) Motion Sequence Generation
Figure 6: Two phase of a collision-prone SE(2) motion sequence is refined. a) A collision occurs on the original path (red T-shape) at a point $p$. The point is then steered to a free position (orange point) which is then connected to existing ones (large green points). b) Child segments with remaining collisions undergo repeated adjustment until achieving a fully collision-free SE(2) motion sequence (purple line).
@@ VI-A Trajectory Generation Strategy
Figure 7: Two refined SE(2) motion sequences for a T-shaped robot navigating an obstacle-dense zone are visualized. (a) Robot kernel collides with obstacles in all $gaps$. (b) A complete, collision-free trajectory composed of two successfully optimized SE(2) segments connected by three $\mathbb{R}^{2}$ segments (orange curve). The successfully SE(2) optimized trajectory pieces in $gap_{3}$ and $gap_{4}$ are shown in green color, while failed in $gap_{2}$ indicates with red.
@@ VI-B Trajectory Formulation
In this work, we adopt MINCO [17], a piece-wise polynomials representation of the trajectory:
$\mathbf{p}(t)=\begin{cases}\mathbf{p}_{1}(t)=\mathbf{c}_{1}^{T}\beta(t-0)&0\leq t<T_{1}\\ \vdots&\vdots\\ \mathbf{p}_{M}(t)=\mathbf{c}_{M}^{T}\beta(t-T_{M-1})&T_{M-1}\leq t<T_{M}\end{cases}$ (2)
Here, the trajectory $p(t)$ is an $m$-dimensional polynomial composed of $M$ segments, each with a degree $N=2s-1$, where $s$ represents the order of the corresponding integrator chain. $c_{i}\in\mathbb{R}^{(N+1)\times m}$ denotes the coefficient matrix of ${M}$ pieces, $\beta(t)=[1,t,\ldots,t^{N}]^{T}$ represents the natural basis,and $T_{i}=t_{i}-t_{i-1}$ is the time duration of the $i$-th segment. The robot’s state can be easily obtained by taking the derivative of the polynomial trajectory. In this paper, we use the degree $N=5$ to ensure the continuity up to snap in adjacent pieces.
@@
$G_{p}(t)=L_{\mu}\left[\|p(t)-p_{i}(t)\|^{2}\right],$ (5)
where $\|\cdot\|^{2}$ denotes the square of the Euclidean norm of a vector. The function $i(t)$ maps to the index of the discretized key node based on the SE(2) result.The pose residual $G_{R}(t)$ is defined as:
$G_{R}(t)=L_{\mu}\left[\|R(t)^{-1}R_{i}(t)-I\|^{2}_{F}\right].$ (6)
Here, $\|A\|^{2}_{F}$ represents the Frobenius norm of matrix $A$, which can be expressed as $\text{tr}(A^{T}A)$ using the matrix trace. Essentially, $G_{R}(t)$ quantifies pose similarity residuals, and $G_{p}(t)$ evaluates position similarity residuals. These optimization components are designed to ensure the closeness between optimization result trajectory and front-end generated motion sequences.

### v2

@@ I INTRODUCTION
Safe navigation of arbitrary-shaped ground robots in constrained environments is an indispensable requirement for real-world autonomy in many scenarios. For example, forklift trucks transporting long timber in the warehouse demand precise whole-body motion planning to avoid collisions in confined aisles. Traditional methods [corridor-polyhedron1, corridor-sphere1, corridor-rectangle1] that conservatively approximate robots with convex shapes largely waste the navigable space, rendering them impractical for such applications. Thus, trajectory planners that explicitly model the robot’s true geometry are necessary for efficient and safe deployment in spatially constrained scenes.
Recent frameworks like Robot-Centric Euclid Signed Distance Field (RC-ESDF) [geng] and Implicit Swept Volume Signed Distance Field (SVSDF) [Zhang_Wang_Xu_Gao_Gao_2023] enable precise geometry-aware robot collision assessment, overcoming the limitations of direct convex approximations. However, RC-ESDF adopts a naive A* searcher that renders a single path candidate, which ignores the possibility of other potentially feasible paths. Furthermore, its discrete sampling collision evaluation in back-end optimization risks missing the sub-resolution narrow collisions in the scene. Unlike RC-ESDF, SVSDF ensures provably accurate Continuous Collision Avoidance (CCA) [SVSDF] by leveraging an iterative swept volume approach. However, the SVSDF planner also employs an A*-based searcher, which suffers the same problem as RC-ESDF’s front end. Moreover, SVSDF’s prohibitive runtime makes it impractical for real-world autonomous applications, which primarily stems from its monolithic back-end formulation that attempts to solve the entire whole-body robot trajectory from start to goal in a single step. In practical ground robot navigation, we observe that finding feasible navigation options rarely succeeds with a single front-end path candidate. Besides, optimizing the whole body SE(2) trajectory is necessary only in localized regions with restricted obstacles, while simpler $\mathbb{R}^{2}$ planning suffices elsewhere. However, both RC-ESDF and SVSDF lack a geometry-aware front end that can provide multiple navigable options, identify critical regions, and adaptively activate whole-body optimization only where needed—leading to computational inefficiency.
@@ II RELATED WORKS
In robot motion planning, existing works [bubble-planner], [kong_drone_sphere] usually model the robot’s geometry as a mass point and inflate the obstacle based on its radius to construct configuration space for collision avoidance. Liu et al. [Liu_Mohta_Atanasov_Kumar_2018] encodes the robot shape as an ellipsoid to search the safe path for drones. Euclidean Signed Distance Fields (ESDF) and corridor-based methods are widely employed to construct collision-aware cost functions for whole-body motion planning trajectory optimization.
@@
The ESDF serves as a map representation that encodes the distance to the nearest obstacle surface. In this approach, the robot’s geometry is often approximated as a sphere [ego_planner], [fast_planner] to simplify distance queries. Collision-free motion is guaranteed by ensuring the ESDF value (i.e., the minimum distance to obstacles) at each trajectory sample point exceeds the radius of the corresponding sphere. While computationally efficient, this spherical approximation may lead to overly conservative trajectories, as the spheres only coarsely enclose the robot’s true geometry. In contrast, corridor-based methods construct trajectories by defining collision-free regions as convex geometric primitives, including polyhedra [corridor-polyhedron1], spheres [corridor-sphere1], and axis-aligned bounding boxes [corridor-rectangle1]. Li et al. [dual-sphere-parking] employs multiple spheres to approximate the robot’s shape, enabling a more flexible and accurate representation of its geometry within the collision-free corridors. These primitives impose linear constraints on the trajectory optimization problem, ensuring the robot’s motion remains within the corridor. To account for the robot’s spatial footprint, the corridor framework often models the robot as an ellipsoid [yunfan-wholebody], a union of spheres [dual-sphere-parking], [Li_Ouyang_Zhang_Acarman_Kong_Shao_2021] and enforces its containment within the corridor’s convex polyhedra. All the above-mentioned methods share a core principle: generating a convex polytope that encapsulates the robot’s geometry to guarantee collision avoidance. However, ensuring continuous collision-free trajectories while preserving the accurate representation of a robot’s non-convex geometric shape remains a significant challenge.
@@
Inspired by the flexibility of the ESDF in representing arbitrary obstacle geometries, Geng et al. [geng] proposed a novel approach that inverts the traditional ESDF paradigm: instead of computing distances from obstacles, the method evaluates distances from the robot’s surface to obstacles. By reformulating the ESDF around the robot’s geometry, the framework actively “pushes” the robot away from obstacles during trajectory optimization. However, this method faces two critical limitations. First, like traditional ESDF, the robot-centric ESDF relies on discrete sampling of the robot’s surface along trajectory which may results in “tunnel effect”. Second, the gradient direction derived from the robot-centric ESDF does not reflect the optimal escape direction for the robot.
@@
To address these issues, SVSDF[SVSDF] innovatively integrates the concept of swept volume into trajectory optimization. It represents the robot’s surface as the zero-level set of an implicit SDF and integrates swept volume to model the space-time continuum of its motion to achieve resolution-independent CCA. Additionally, the SVSDF is calculated by finding the radius of the smallest sphere at an obstacle point that is tangent to the swept volume boundary which is always along the optimal gradient for collision avoidaance. However, precisely considering the robot’s actual geometry in a single optimization problem from start to end for the continuous collision-free trajectory is time-consuming. It causes much pressure for the optimizer to converge to a satisfied minimum. Thus, we present an adaptive coarse-to-fine trajectory generation paradigm, accurately representing the robot’s geometry only when it is needed, and using simpler planners elsewhere to speed up the process.
@@ IV Topological Path Generation
To maximize the exploration of various topological paths in a narrow space while maintaining efficiency, in this stage, we use the minimum inflation margin which is the inscribed circumference of the robot as shown in Fig. 3(a), to inflate the occupancy grid map, representing the coarse navigability at this initial stage. After that, we construct topological roadmaps based on [topology] with a rough consideration of the robot’s shape.
@@ 6    $p_{new}\leftarrow\neg\textbf{Safe}$;
7    while $\neg\textbf{Safe}(p_{new})$ $\land$ $N_{attempt}$ << $N_{max}$ do
@@ 11       $\mathcal{M}_{yaw}\leftarrow\textbf{UpdateState}(\mathcal{\theta}_{new},p_{new},\mathcal{M})$;
In [topology], redundant zigzag paths in obstacle-dense regions are shortened while treating the robot as a mass point and ignoring the actual shape. This will result in missing regions that can be passed in a particular orientation. Therefore, we redesigned the shortcut process to generate a refined topologically equivalent shortcut path $P_{s}$ considering the any-shape robot’s physical geometry (refer to Alg. 1 and Fig. 4). Initially, we denote the original detoured path (blue line) found by the depth-first search from the roadmap as $P_{i}$ and the expected final simplified path (green line) as $P_{s}$. Then, the algorithm uniformly discretizes $P_{i}$ into a set of points $P_{d}=\{p_{1},p_{2},...,p_{n}\}$ (blue point in the path) based on the grid map resolution and initializes the simplified path $P_{s}$ with the start point $p_{1}$ (Line 1). For each point $p_{d}$ in $P_{d}$, the algorithm performs visibility checking between $P_{s}.back()$ and $p_{d}$ as in [topology] (Line 2). When visibility is blocked by an obstacle, the algorithm identifies the obstruction point as $p_{c}$ (Line 4).
@@
Then, we expect to push the obstructed configuration in $p_{c}$ away from the occupancy to generate new collision-free waypoints. Inspired by [geng], to handle the arbitrary-shaped robot, we construct body-frame ESDF inside the robot geometry $\mathcal{M}$ centered at obstruction point $p_{c}$ to render the obstacle-avoiding direction, as shown in Fig. 3(b). The robot’s initial orientation at the obstruction point $p_{c}$ (red T-shape) is important to construct the appropriate body-frame ESDF for guiding collision point adjustment. It is determined by applying Cubic Spline Interpolation between the initial and final orientation at $P_{s}.back()$ and $p_{d}$ (blue T-shape), respectively, and then storing the resulting orientation with the geometric representation and position in $\mathcal{M}_{yaw}$ for generating the ESDF (Line 5). we build an Axis-Aligned Bounding Box (AABB) to extract all nearby obstacle points $X_{obs}$ relative to $p_{c}$ (Line 8). The SDF value at any position $x_{obs}$ in $X_{obs}$ is defined as the minimum distance from it to the nearest robot surface, with positive values indicating the robot’s exterior and negative values indicating its interior, which can be efficiently computed via a generalized winding number method [generalizedwindingnumber] within LIBIGL [^2] https://libigl.github.io/ (Line 9). The gradient in the robot local frame $\nabla\mathcal{SDF}_{robot}^{\mathcal{M}}$ is defined as the direction of the steepest ascent to the nearest robot surface from the query point $x_{obs}$ and computed using central finite differences of distance values between adjacent grid cells. Subsequently, the gradient of the robot SDF in the world frame $\nabla\mathcal{SDF}_{world}^{\mathcal{M}}$ which directs away from the obstacle is obtained with a proper coordinate transform as follows:
@@
where ${R}^{-1}(p_{c})$ is the inverse of the robot’s rotation matrix at $p_{c}$. The norm and direction of the gradient (pink arrow) can push the center of robot ESDF in obstruction point $p_{c}$ away from the obstacle with a magnitude based on the SDF value and minimum safe margin (Line 10). The resultant position after applying SDF-based steering once might remain collided with the confined environments. Thus, we repeat this process within the maximum allowed attempts (Line 7). If the robot collides after reaching the limitation attempts, the last $p_{new}$ along with its orientation $\theta_{new}$ is temporally designated as a new waypoint for further SE(2) refinement (Line 12). The refining process continues until the last point is reached. After that, uniform visibility deformation (UVD) [topology] is implemented to distinguish topological equivalence and eliminate redundant paths effectively. Finally, Alg. 1 generates a new shortened and smoother path consisting of a series of waypoints with feasible orientation stored according to the actual geometry of the robot while preserving the topology attribute.
@@ V SE(2) Motion Sequence Generation
Figure 6: Two phase of a collision-prone SE(2) motion sequence is refined. a) A collision occurs on the original path (red T-shape) at a point pp. The point is then steered to a free position (orange point) which is then connected to existing ones (large green points). b) Child segments with remaining collisions undergo repeated adjustment until achieving a fully collision-free SE(2) motion sequence (purple line).
@@ VI-A Trajectory Generation Strategy
Figure 7: Two refined SE(2) motion sequences for a T-shaped robot navigating an obstacle-dense zone are visualized. (a) Robot kernel collides with obstacles in all gapsgaps. (b) A complete, collision-free trajectory composed of two successfully optimized SE(2) segments connected by three $\mathbb{R}^{2}$ segments (orange curve). The successfully SE(2) optimized trajectory pieces in $gap_{3}$ and $gap_{4}$ are shown in green color, while failed in $gap_{2}$ indicates with red.
@@ VI-B Trajectory Formulation
In this work, we adopt MINCO [MINCO], a piece-wise polynomials representation of the trajectory:
$\mathbf{p}(t)\!=\!\begin{cases}\mathbf{p}_{1}(t)=\mathbf{c}_{1}^{T}\beta\left(t-0\right)&0\leq t<T_{1}\\ \vdots&\vdots\\ \mathbf{p}_{M}(t)\!=\!\mathbf{c}_{M}^{T}\beta\left(t-T_{\!M\!-\!1}\right)\!\!\!&T_{\!M\!-\!1}\!\leq\!t\!<\!T_{\!M}\end{cases}$ (2)
Here, the trajectory $p(t)$ is an $m$-dimensional polynomial composed of $M$ segments, each with a degree N=2s1N=2s-1, where $s$ represents the order of the corresponding integrator chain. $c_{i}\in\mathbb{R}^{(N+1)\times m}$ denotes the coefficient matrix of M{M} pieces, $\beta(t)=[1,t,\ldots,t^{N}]^{T}$ represents the natural basis,and $T_{i}=t_{i}-t_{i-1}$ is the time duration of the $i$-th segment. The robot’s state can be easily obtained by taking the derivative of the polynomial trajectory. In this paper, we use the degree $N=5$ to ensure the continuity up to snap in adjacent pieces.
@@
$G_{p}(t)=L_{\mu}\left[\|p(t)-p^{i(t)}\|^{2}\right],$ (5)
where $\|\cdot\|^{2}$ denotes the square of the Euclidean norm of a vector. The function $i(t)$ maps to the index of the discretized key node based on the SE(2) motion sequence. The rotation residual $G_{R}(t)$ is defined as:
$G_{R}(t)=L_{\mu}\left[\|R(t)^{-1}R^{i(t)}-I\|^{2}_{F}\right].$ (6)
Here, $\|A\|^{2}_{F}$ represents the Frobenius norm of matrix $A$, which can be expressed as $\text{tr}(A^{T}A)$ using the matrix trace. Essentially, $G_{R}(t)$ quantifies rotation similarity residuals, and $G_{p}(t)$ evaluates position similarity residuals. These optimization components are designed to ensure the closeness between optimization result trajectory and motion sequences.